OpenStudy (anonymous):

Write the equation of the horizontal line passing through the point (-1, -8).

OpenStudy (anonymous):

It's a horizontal line, so the value of y stays the same (-8).

OpenStudy (anonymous):

can you write an equation for a line that stays on -8 no matter the value of x?

OpenStudy (anonymous):

um an equation as in?

OpenStudy (anonymous):

y = n * x, or y = n+x or something like that

OpenStudy (anonymous):

For example y=2x would look like